(TCO 5) Which of the following is true?

Group of answer choices

Current yield = dividends / price paid.

Coupon rate = interest / price paid.

YTM = interest / 30.

None of the above

(TCO 8) Who would not normally be concerned about creating an investment policy for portfolio creation?

CEO of a firm issuing bonds

401k plan manager

Pension fund manager

None of the above

(TCO 4) Which of the following would be a good indicator of the bond market?

Group of answer choices

Lehman Brothers Index

S&P 500

CBOE

None of the above

(TCO 6) The yield on a corporate bond is 10%, and it is currently selling at par. The marginal tax rate is 20%. A par value municipal bond with a coupon rate of 8.50% is available. Which security is a better buy?

Group of answer choices

Municipal bond

Both are equal

Corporate bond

A municipal bond carries no par
Answer #1

1. The answer is "Coupon rate = Interest / Price paid"

suppose a 1000 $ bond paying 20 $ as coupon payment annually. Therefore coupon rate = interest / bond value * 100

= 20 / 1000 * 100 = 2%

2. The answer is "CEO of a firm issuing bond"

CEO is responsible for management of business enterprise rather than creating investment portfolio.

3.The answer is "Lehman brothers index"

A benchmark index that tracks Government/Corporate Bond market.

4. The answer is "Municipal Bond"

Municipal Bond return = .8.50 % (tax free return)

corporate Bond return = 10 %

Marginal Tax rate = 8.50 %

So after tax return of corporate bond = before tax return * (1 - Marginal tax rate)

= 10% * (1 - 20%)

= 10 * 0.80

= 8%

Therefore municipal bond return is greater by 0.50%.